# CONLINE: Complex Code Generation and Refinement with Online Searching and Correctness Testing

### v1

@@ 4.3 Ablation Study
In this work, to validate the effectiveness of CONLINE, we conduct different ablation studies, with results presented in Table 1 and Table 2. Details on the ablation study are shown in §C.
CONLINE significantly enhances the base model’s ability to generate complex code. Compared to the base model, CONLINE (GPT-4) has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CONLINE.
Online search, generation of test cases, and serialization each contribute to the model’s performance improvements. Compared to CONLINE, after performing ablation studies, these features showed a decrease in performance of 3.9%, 5.9%, and 1.3% respectively on the DS-1000 dataset. The online search improves the model by providing common code patterns, which reduces the difficulty of the model in generating initial code. Serialization, by converting inputs and outputs into a sequential format, allows the model to more intuitively observe inputs and outputs that are too lengthy or are in non-textual modalities, thereby strengthening its ability to solve complex code problems.
During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important. After separately conducting ablation studies on the output result and error, CONLINE shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval datset. This indicates that the output result contributes more to the refinement process than the error. However, in previous works, the output result is often overlooked, which should not be the case, especially in the generation of complex code. The evidence from the ablation study emphasizes the necessity of paying attention to the output results during the refinement phase to ensure the generation of high-quality, complex code.
@@ 4.4 Analysis of Different Base Models Performance
Table 3: Table of Different Base Models Results for DS-1000. All metric numbers are represented as percentages. For each metric in each section, the bold number indicates the highest performance.
Method Perturbation Total/Avg.
Origin Surface Semantic Diff-Rewrite
GPT-4 64.47 69.74 56.96 43.83 60.20
   w/ retrieve 66.23 71.05 59.57 45.68 62.10
   w/ refine 68.64 70.39 60.00 51.23 64.10
   CONLINE 71.71 74.34 66.96 53.09 68.00
GPT-3.5 57.02 43.42 40.00 32.72 47.10
   w/ retrieve 47.15 25.00 36.96 25.31 37.90
   w/ refine 55.70 50.66 44.35 35.80 49.10
WizardCoder 41.01 21.71 31.74 16.05 31.90
   w/ retrieve 15.79 9.21 12.17 9.88 13.00
   w/ refine 39.69 21.71 30.00 15.43 30.80
Table 3 shows the performance results of CONLINE on the DS-1000 dataset with different base models. We can see that GPT-4 has been comprehensively improved with CONLINE, but the performance on GPT-3.5 and WizardCoder is mixed. This indicates that CONLINE requires the model to have the following capabilities to enhance its performance:
For code generation planning ability, the higher the complexity of the code that needs to be generated, the higher the demand for planning ability. Planning capability is key to online retrieval; only correct planning can generate appropriate queries to retrieve useful information. After incorporating online retrieval, GPT-3.5 has an increase of 4.75% on ClassEval as shown in Table 2, yet it decreased by 9.2% on DS-1000 as shown in Table 3. For ClassEval, we plan for each function during online retrieval. The challenge of ClassEval lies in how to generate the entire class and the interrelated functions, but the complexity of individual function codes is not as high as that of DS-1000. Therefore, the planning ability of GPT-3.5 can handle ClassEval, but it is inferior on DS-1000.
Code generation also necessitates the models to possess the capability of in-context learning. The generated code should be built on all the above provided contents, and the understanding of the generated results based on the preceding input prompt is of great importance in the refinement stage. In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively. And the overall performance of WizardCoder on DS-1000 is comparatively interior to GPT models. One possible reason for this lies in the limited in-context learning capability of WizardCoder given some of prompts of DS-1000 are lengthy and complicated, which makes it difficult for WizardCoder to comprehend the related context and make corresponding code modifications accurately.
4.5 Case Study
Figure 3: Case Study for Correctness Testing.
@@ 4.5 Case Study
For the case study on correctness testing, refer to Figure 3. It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results. The model refines the code based on the output, thereby improving subtle errors and generating the correct code.
@@ 5 Conclusion
In conclusion, CONLINE introduces a novel framework for generating complex code in real-word scenarios by emulating human coding processes like online searching and test case creation. It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data. The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding. Tested on various datasets, CONLINE outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
@@
Appendix A Details for on-line search
@@
Appendix B Details for Base Models
@@
Appendix C Details for Ablation Study
@@
Appendix D Case Study
Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset. Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices. After unsuccessful attempts, the model turns to an online search through a self-generated query and integrating the research results as part of next prompt. Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class. Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
@@ Appendix D Case Study
Appendix E Prompts of CONLINE

### v2

@@ 4.3 Ablation Study
Table 3: Different Base Models Results for DS-1000 and ClassEval. All metric numbers are represented as percentages. For each metric in each section, the bold number indicates the highest performance.
Method DS-1000 ClassEval
Origin Surface Semantic Diff-Rewrite Total/Avg. Class-level Method-level
GPT-4 64.47 69.74 56.96 43.83 60.20 43.5 67.3
     + retrieve 66.23 71.05 59.57 45.68 62.10 50.6 72.4
     + refine 68.64 70.39 60.00 51.23 64.10 44.9 69.8
     CoCoST 71.71 74.34 66.96 53.09 68.00 52.8 77.6
GPT-3.5 57.02 43.42 40.00 32.72 47.10 35.4 59.4
     + retrieve 47.15 25.00 36.96 25.31 37.90 41.9 61.7
     + refine 55.70 50.66 44.35 35.80 49.10 42.8 62.3
     CoCoST - - - - - 45.8 64.7
WizardCoder 41.01 21.71 31.74 16.05 31.90 23.0 51.1
     + retrieve 15.79 9.21 12.17 9.88 13.00 18.2 41.8
     + refine 39.69 21.71 30.00 15.43 30.80 22.3 50.7
In this work, to validate the effectiveness of CoCoST, we conduct different ablation studies, with results presented in Table 1 and 2. Details on the ablation study are shown in §B.4.
CoCoST significantly enhances the base model’s ability to generate complex code. Compared to the base model, CoCoST has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CoCoST.
Online search, generation of test cases, and serialization each contribute to the model’s performance improvements. Compared to CoCoST, after performing ablation studies, these features showed a decrease in performance of 3.9%, 5.9%, and 1.3% respectively on the DS-1000 dataset. The online search improves the model by providing common code patterns, which reduces the difficulty of the model in generating initial code. Serialization, by converting inputs and outputs into a sequential format, allows the model to more intuitively observe inputs and outputs that are too lengthy or are in non-textual modalities, thereby strengthening its ability to solve complex code problems.
Online search outperforms offline retrieval in effectiveness and has a wider range of applicability. As shown in Table 1, using only online retrieval (the row w/o generation of test case) outperforms DocPrompting, which is an offline retrieval approach. Moreover, in real-world scenarios, as opposed to specific datasets, the types of problems encountered are more diverse. The scalability of online retrieval enables them to effectively address a wide range of problems. However, offline retrieval systems struggle to encompass all relevant information comprehensively.
During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important. After separately conducting ablation studies on the output result and error, CoCoST shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval dataset. This indicates that the output result contributes more to the refinement process than the error. However, in previous works, the output result is often overlooked, which should not be the case, especially in the generation of complex code. The evidence from the ablation study emphasizes the necessity of paying attention to the output results during the refinement phase to ensure the generation of high-quality, complex code.
@@ 4.4 Analysis of Different Base Models Performance
Table 3 shows the performance results of CoCoST on the DS-1000 dataset with different base models. We can see that GPT-4 has been comprehensively improved with CoCoST, but the performance on GPT-3.5 and WizardCoder is mixed. This indicates that CoCoST requires the model to have the following capabilities to enhance its performance:
For code generation planning ability, the higher the complexity of the code that needs to be generated, the higher the demand for planning ability. Planning capability is key to online retrieval; only correct planning can generate appropriate queries to retrieve useful information. After incorporating online retrieval, GPT-3.5 has an increase of 4.75% on ClassEval, yet it decreased by 9.2% on DS-1000 as shown in Table 3. The challenge of ClassEval lies in how to generate the entire class and the interrelated functions, but the complexity of individual function codes is not as high as DS-1000. Thus, the planning ability of GPT-3.5 can handle ClassEval, but it is inferior on DS-1000.
Code generation necessitates models to have in-context learning abilities. The generated code should be built on all the above provided contents, and the understanding of the preceding input prompt is of great importance in the refinement stage. In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively. And the overall performance of WizardCoder is comparatively interior to GPT models. This could be due to WizardCoder’s limited in-context learning ability, especially with the complex and lengthy prompts, hindering accurate context comprehension and code modification.
4.5 Case Study and Error Analysis
@@
For the case study on correctness testing, refer to Figure 3. It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results. The model refines the code based on the output, thus improving hidden errors and generating the correct code.
Our framework consists of multiple components cascaded together, which results in certain intermediate steps that cannot be explicitly validated for effectiveness, as well as the potential generation of cascading errors. For the former, a discussion is provided in §C.2, while for the latter, an error analysis is conducted in §C.3.
@@ 5 Conclusion
In conclusion, CoCoST introduces a novel framework for generating complex code in real-world scenarios by emulating human coding processes like online searching and test case creation. It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data. The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding. Tested on various datasets, CoCoST outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
@@
Appendix A Online Searching Detail
Website Base Station Distributions Table During the DS-1000 Online Retrieval Process:
@@ Appendix A Online Searching Detail
Moreover, CoCoST can be applied to specialized, proprietary, or domain-specific knowledge repositories as long as they are accessible via query. Moreover, implementing queries for private datasets is easily achievable and a growing trend in data management. Major companies such as Google and Microsoft already offer products designed to search private data; for example, Google Workspace’s Cloud Search provides powerful capabilities for enterprises to search their private data. In this paper, to validate the effectiveness of our framework, we conducted tests on public online searches. Moving forward, the framework can be applied to an even broader range of knowledge repositories.
Appendix B Experiment
B.1 Datasets Detail
Further details of DS-1000 implementation are as follows:
  • We implement a filtering approach to prevent data leakage and model replication of existing solutions from Stack Overflow. The DS-1000 dataset originates from Stack Overflow, and concurrently, over 50% of the websites we encountered during our online searches are from Stack Overflow. Thus, to prevent data leakage, when conducting online searching, we filter out all Stack Overflow problems belonging to the source of the DS-1000 dataset by using the Stack Overflow question_id.
B.2 Base Models
@@
B.3 Baselines Details
  • DocPrompt Zhou et al. (2023): DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools. The model then conditions on these documents, along with the problem description, to generate code.
  • Self-Debugging Chen et al. (2023): This approach depends on a SQL application or Python interpreter to instruct language models in revising SQL commands or Python code containing errors. For the sake of a fair comparison, we utilize its "simple" variant.
  • SELFEVOLVE Jiang et al. (2023): Employs LLMs as both sources of knowledge and self-reflective programmers. During the self-reflective process, it refines the code by addressing bugs.
  • Reflexion Shinn et al. (2023): Reflexion utilize reflective feedback with generated tests and episodic memory to process task feedback. For the sake of a fair comparison, we utilize GPT-4 as base model and set trail number = 1.
It is worth noting that the test cases involving the refinement process in the baselines mentioned above all use the test cases from the dataset designated for testing. However, within the context of the real-world scenario of CoCoST, test cases from the dataset should not be used within the framework. Without these test cases, they are entirely incapable of functioning.
B.4 Ablation Study Details
@@
    Appendix C Experimantal Results
    C.1 DS-1000 Results
    The main results for different packages in DS-1000 are shown in Table 5.
    Table 5: Table of Main Results for different packages in DS-1000. All metric numbers are represented as percentages. The bold number indicates the highest performance.
    Method Pandas Numpy Matplotlib Tensorflow Scipy Sklearn Pytorch Total/Avg.
    CoCoST 59.45 75.91 75.48 71.11 61.32 63.48 77.94 68.00
         + retrieve 51.89 70.91 68.39 66.67 52.83 70.43 60.29 62.10
         + refine 55.67 72.73 74.19 64.44 54.72 60.00 70.59 64.10
         GPT-4 only 52.23 70.45 67.74 55.56 50.00 64.35 55.88 60.20
    The results indicate that CoCoST shows a more pronounced effect on libraries whose inputs and outputs are more complex or more challenging for LLMs to intuitively understand, such as Matplotlib, TensorFlow, and PyTorch. On Sklearn, CoCoST experiences a slight decline due to its test cases containing complex objects, which present a significant challenge in generating test cases. Consequently, CoCoST’s performance on Sklearn is not as strong as with the other libraries.
    C.2 Analysis of Pipeline
    Regarding the generation of test cases, to prove that the generated test cases are comparable to the ground truth test cases, we substitute the generated test cases in CoCoST with ground truth test cases and conduct experiment on the DS-1000 dataset using GPT-4. The results showed that the performance was 68.70%, only marginally higher by 0.7% compared to the use of generated test cases (68.00%). This proves that the effects of both are comparable and that the generated test cases rarely lead to errors.
    Regarding generating plans, we attempt to bypass the planning step and directly perform online retrieval. On the DS-1000 dataset using GPT-4, this approach result in a performance of 55.70%, which is 6.4% lower than using planning for online retrieval (62.10%), and even 4.5% lower than only using GPT-4 to generate code (60.20%). This confirms that the generated plans are significantly effective.
    C.3 Error Analysis
    For cascading errors, some errors generated by retrieval processes can be corrected through refinement, while others may persist, necessitating future improvements. Take DS-1000 as an example: Compared to the baseline (GPT-4), we observed that among the instances that turned erroneous after retrieval, 39.6% were corrected, while 60.4% remained incorrect. For the former cases, it shows that CoCoST is able to fix some bad cases in the refinement stage even though the retrieval contents have some errors. For the later cases, we do observe some cases that are worthy of further research as potential directions for future work. First, the search content could be more detailed. E.g., Some basic steps that LLMs consider unnecessary to search for are not generating queries, but is exactly where the bug in the code. Second, the search query could be more targeted. The descriptions of some queries are not specific enough in terms of some complicated problems. Therefore, more sub-queries are needed to help model receive clearer instructions.
C.4 Case Study
Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset. Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices. After unsuccessful attempts, the model turns to an online search through a self-generated query and integrates the research results as part of the next prompt. Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class. Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
Figure 3: Case Study for Correctness Testing.
@@ C.4 Case Study
Appendix D Prompts of CoCoST